OpenStudy (anonymous):

Factor: z2 - 2z A. z(z + 2) B. (z - 1)(z - 2) C. z2 D. z(z - 2)

OpenStudy (anonymous):

well try it by your self its easy as pie

OpenStudy (anonymous):

How do I do it? Could you give me steps?

OpenStudy (anonymous):

Is it C?

OpenStudy (anonymous):

yup good job

OpenStudy (anonymous):

\[z^2 - 2z\] \[z(z - 2)\]

OpenStudy (anonymous):

since you can factor out a z from both terms. \[z \times z = z^2\] \[x \times -2 = -2x\]

OpenStudy (anonymous):

@AntoinetteWillman & @prettygirlbabyblue15 how did you guys get C?

OpenStudy (anonymous):

oh sorry didnt pay attention

OpenStudy (anonymous):

to that one

OpenStudy (anonymous):

can you help me on my question @some_someone

OpenStudy (anonymous):

@AntoinetteWillman do you understand?

OpenStudy (anonymous):

idk what is your question?

OpenStudy (anonymous):

@prettygirlbabyblue15

OpenStudy (anonymous):

Jamie and Nicole are buying streamers for their mother’s surprise birthday party. They need to buy purple and gold streamers. The purple streamers cost $2 per roll and the gold streamers cost $3 per roll. They have $25 to spend on streamers. Which of the following is a possible combination of streamers? A. 1 roll of purple streamer and 8 rolls of gold streamer B. 3 rolls of purple streamer and 6 rolls of gold streamer C. 7 rolls of purple streamer and 4 rolls of gold streamer D. 9 rolls of purple streamer and 3 rolls of gold streamer

OpenStudy (anonymous):

@some_someone

OpenStudy (anonymous):

i think its B @some_someone